# Deontic Temporal Logic for Formal Verification of AI Ethics

### v2

@@ I Introduction
To address this challenge, we are exploring the use of formal methods to express and prove the ethical correctness of AI systems. One promising approach is the use of deontic logic, a branch of modal logic that deals with concepts such as obligation, permission, and prohibition [23, 24]. Deontic logic provides a rigorous framework for reasoning about ethical norms and can be used to formalize ethical principles [25] and constraints. Several works have explored using deontic logic to formalize machine ethics mainly for robots [26, 27] and normative systems [28]. These studies have concentrated on Kantian ethics, integrating deontic and temporal logic to verify the ethical behavior of autonomous systems, such as unmanned aircraft, over time [29, 30].
@@
While promising, these methods are often constrained by specific ethical frameworks and fail to scale with modern AI’s complexity, which increasingly mimics human tasks, leverages natural language processing, and operates on vast datasets. This leads to a proliferation of potentially subjective ethical rules influenced by personal biases. The dynamic, evolving nature of AI further complicates ethical formalizations [31]. Critically, many of these approaches remain theoretical, lacking practical integration with machine learning techniques, highlighting the need for more adaptive and implementable ethical frameworks in AI [32]. Our work represents a foundational effort to develop a unified framework that addresses ethical principles in AI systems, specifically focusing on granular levels of explainability and fairness. It builds upon existing approaches [33, 26] in specialized domains, extending them to tackle the unique challenges posed by modern AI ethics. In doing so, this paper introduces a novel direction for formalizing and verifying the ethical principles of AI systems.
@@
While the proposed framework is broadly applicable to the ethics verification of autonomous systems, this work specifically focuses on its application to AI. The scope of this work is to provide a conceptual foundation and framework for formalizing AI ethics using deontic and temporal logic. Rather than focusing on individual actions or decisions, our approach emphasizes system-level specifications. It involves defining ethical properties that AI systems should ideally meet. For example, an AI system that uses gender as a feature should avoid making decisions explicitly based on it. These properties ensure that systems are designed to identify and mitigate biases and ethical violations effectively. Defining properties like "forbidden to consider sensitive features in predictions," provides a way for analysis of commonly discussed properties of AI in a unified manner. This abstraction also helps to reduce the difficulty in formalizing each action or each type of user [31]. Additional details on formalizing the properties of AI systems and their verification are provided in Section IV.
@@
The basic model for applying deontic logic to AI ethics uses first-order logic to define predicates (Table I) and axioms that capture ethical requirements. This model introduces variables and predicates such as $\mathit{x}$ to indicate an AI system, $\mathit{x}_{a}$ to indicate that system $\mathit{x}$ performs an action $\mathit{a}$, and $\mathcal{E}(x)$ to indicate ethical behavior. Axioms 1.1,  1.2,  1.3, and  1.4 are then defined using these predicates to express ethical obligations, prohibitions, and permissions for AI systems. Building upon the basic model, an extended version incorporates temporal operators from temporal logic to reason about the ethical behavior of AI systems over time. This extension allows for the expression of more complex ethical requirements, such as the obligation for AI systems to maintain fairness over time or the prohibition of exhibiting bias. The temporal operators used in this model include “always” ($\Box$), “eventually” ($\diamond$), and “until” ($\mathcal{U}$) presented by Manna and Pnueli [34]. In this framework, $\mathit{x}$ persists across states, while actions cause state transitions. If no action influences the transition, the passage of time follows a default evolution independent of AI choices.
@@
Theorems III.2 and III.3 in the basic model of deontic logic for AI ethics explore the relationships between ethical obligations, prohibitions, and permissions for AI systems. These theorems employ first-order logic and the defined predicates to derive conclusions about the ethical behavior of AI systems. The proofs of these theorems rely on techniques such as modus ponens, contraposition, and proof by contradiction to establish the logical connections between the axioms and the derived statements. The general flavor of the theorems is to provide a rigorous foundation for reasoning about the ethical requirements of AI systems, demonstrating how the Axioms 1.11.21.3, and  1.4 can be used to infer specific obligations, prohibitions, and permissions in various contexts. By establishing these logical relationships, the theorems contribute to a comprehensive framework for analyzing and ensuring the ethical behavior of AI systems. Similar to this, the Theorems III.5 to III.14 in the extended model, which incorporates temporal logic operators, explore the ethical behavior of AI systems over time. These theorems focus on capturing the temporal aspects of ethical requirements, such as the obligation to maintain fairness or the prohibition of exhibiting bias. The proofs of these theorems utilize the semantics of the temporal operators, such as $\Box$(always), $\diamond$(eventually), and $\mathcal{U}$(until), in conjunction with the Axiom lists 2 to 3 and predicates defined (Table I) in the basic model. The general flavor of the theorems in the extended model is to provide a more expressive and nuanced framework for reasoning about the ethical behavior of AI systems, considering the dynamic and evolving nature of these systems. The theorems establish logical connections between the temporal properties of AI systems and their ethical obligations, allowing for the analysis of more complex and realistic scenarios. By incorporating temporal aspects, the extended model enables a deeper understanding of the long-term ethical implications of AI systems and provides a foundation for designing and verifying AI systems that behave ethically over time.
@@
The importance of this work lies in formalizing the axioms to define an AI system’s ethical requirements and its potential to provide a formal and verifiable framework for ensuring the ethical behavior of AI systems. Our experimental findings show the effectiveness of this formalization in assessing the ethics of real-world AI systems—Loan prediction and COMPAS. We evaluated the ethical aspects of the systems, wherein we defined specific properties that these systems must adhere to in order to be deemed ethical. Our results revealed that certain properties were indeed satisfied by the system, while others were not (Table III). The results demonstrated that applying deontic logic and temporal operators to AI ethics represents a significant step forward to formally specify and verify the ethical behavior of AI systems.
@@ II Related Works
The earlier work introduce the GenEth ethical dilemma analyzer [37], which utilizes inductive logic programming to infer principles for ethical actions. Dominance Act Utilitarianism (DAU), a deontic logic of agency, is another framework for encoding and analyzing obligations in autonomous systems. DAU frameworks are efficient in addressing safety-critical behaviors, such as adherence to traffic laws and avoidance of reckless actions [38]. Such frameworks can formalize ethical obligations in systems like self-driving cars, enabling systematic reasoning about social and moral responsibilities [39]. Additionally, several works employ the Belief-Desire-Intention (BDI) framework to formalize reasoning about moral agents [40, 41, 30]. This structure supports transparency and formal verification in ethical decision-making processes for robots.
Further, the literature explores the use of high-level action languages and Answer Set Programming to design ethical autonomous agents [42, 43, 44]. There are several works that propose using deontic logic to constrain robot behavior in ethically sensitive environments, as this type of logic helps interpret natural language directly [45, 26, 27]. This frameworks are also used for ethical reasoning in the healthcare domain, emphasizing accountability and transparency [46]. Additionally, various works focus on using deontic logic-based frameworks for formalizing ethical reasoning in AI systems [47, 48, 49].
@@
Therefore, rule-based ethical theories, particularly deontology, are essential for developing trustworthy AI systems [53]. However, considering the dynamic nature of AI, especially regarding fairness and explainability at a granular level, significant gaps remain. Our work addresses these gaps by introducing fairness and explainability at multiple granularities, including stable, transient, inherent, and retrofitted/artificial dimensions. These distinctions capture the evolving nature of AI and its complex decision-making processes, providing a more comprehensive approach to ethical verification. Furthermore, existing frameworks often overlook the impact of personal biases introduced during training and lack mechanisms to mitigate them effectively. To address this, we propose an iterative learning approach designed to identify and reduce the influence of personal biases in the system. While prior research highlights the gap between theoretical ethical reasoning and its practical application in autonomous agents [32], our framework bridges this divide. By implementing and testing the framework in real-world AI systems such as COMPAS and Loan prediction systems, we validate its effectiveness and ensure its applicability.
@@ III-A Preliminaries
Standard Deontic Logic (SDL) and Temporal Deontic Logic (TDL) represent two distinct variations within deontic logic. We incorporate both SDL and its extension, TDL, into our work. SDL formulas include classical propositional logic and it operates as a monadic deontic logic, meaning its operators (obligation, permission, forbidden) apply to individual formulas ($\varphi$); they are read as "it is obligatory that $\varphi$", "it is permissible that $\varphi$", and "it is forbidden that $\varphi$" respectively. Furthermore, they are cross-definable. For instance, $P\varphi:=\neg O(\neg\varphi)$, and $F\varphi:=O\neg\varphi$ This logical statement explains that permission ($P$) or forbidden ($F$) can be represented in terms of obligation ($O$). Temporal Deontic Logic expands SDL by integrating temporal aspects into norms and obligations, introducing operators such as $always(\Box),eventaully(\diamond),next$, and $until(\mathcal{U})$. For instance, $next\ \varphi$ means that the proposition $\varphi$ holds in the next time step. Similarly, $\varphi\mathcal{U}\psi$ means $\varphi$ is true until $\psi$ becomes true. We use the semantics of the combined logic as an extension of the Kripke-style possible world semantics of deontic logic with temporal operators, as described in reference [54]. While branching-time logic is often used to reflect future uncertainty, we adopt Linear Temporal Logic (LTL) due to its simplicity and relevance to AI verification, where obligations typically unfold along a single execution path. The alternativeness relation in our model preserves sequential consistency rather than enabling arbitrary time jumps. We recommend interested readers to refer [55, 54] for further details on the foundational principles.
In this section, we focus on formalizing the overall ethical behavior of an AI system. For the formalization, we use the predicates as shown in Table I. The predicate is a function that takes an input and returns a truth value. The following set of axioms has been formulated to articulate the necessary and sufficient characteristics for an AI system to be considered ethical.
@@
Given a set $p$ of atomic proposition, the temporal deontic logic, TDL is defined as,
@@
    If an AI system $\mathit{x}$ is ethical, then it is obligatory that $\mathit{a}$ is an ethical action and that $\mathit{x}$ performs $\mathit{a}$: $\mathcal{E}(x)\rightarrow\mathcal{O}(\mathit{x_{a}}\land\mathcal{E}(\mathit{a}))$.
@@
    An ethical AI system $x$ is forbidden to perform an unethical action $\mathit{a}$: $\mathcal{E}(x)\rightarrow\neg P(\mathit{x_{a}}\land\neg\mathcal{E}(a))$.
@@
    For an ethical AI system $x$, performing an ethical action $\mathit{a}$ is permitted: $\mathcal{E}(x)\rightarrow P(\mathit{x_{a}}\land\mathcal{E}(\mathit{a}))$.
@@
    An AI system following ethical guidelines performs ethical actions: $\mathcal{G}(x)\rightarrow\mathcal{E}(a)$
@@
Based on the foundational principles outlined above and the domain knowledge, we have formulated Theorems  III.2 and  III.3 to ensure that the AI system adheres to ethical standards in all relevant dimensions. Theorem III.2 asserts that for an AI system to be ethical, it must follow all the ethical guidelines. It specifies that the system should not only be accurate in predicting outcomes but also do so in a manner that upholds principles such as fairness, transparency, and non-discrimination.
Theorem III.2.
@@
Assume $O(\mathit{x_{a}}\land\mathcal{E}(\mathit{a}))$ (AI system is obliged to perform ethical action).
[…]
Theorem III.3 states that an AI system cannot be in a state where it is required to do something but not allowed to do it. This maintains logical and ethical consistency. Such a condition is fundamental in designing AI systems that reason ethically, as it ensures they are never blocked from doing what is morally required.
Theorem III.3.
@@
While Theorems  III.2 and  III.3 contribute to formalizing the general ethics of an AI system, it remains essential to develop rigorous formalizations for each individual ethical principle.

### v3

@@ I Introduction
To address this challenge, we are exploring the use of formal methods to express and prove the ethical correctness of AI systems. One promising approach is the use of deontic logic, a branch of modal logic that deals with concepts such as obligation, permission, and prohibition [23, 24]. Deontic logic provides a rigorous framework for reasoning about ethical norms and can be used to formalize ethical principles [25] and constraints. Several works have explored using deontic logic to formalize machine ethics, mainly for robots [26, 27] and normative systems [28]. These studies have concentrated on Kantian ethics, integrating deontic and temporal logic to verify the ethical behavior of autonomous systems, such as unmanned aircraft, over time [29, 30].
@@
While promising, these methods are often constrained by specific ethical frameworks and fail to scale with the complexity of modern AI, which increasingly mimics human tasks, leverages natural language processing, and operates on vast datasets. This leads to a proliferation of potentially subjective ethical rules influenced by personal biases. The dynamic, evolving nature of AI further complicates ethical formalizations [31]. Critically, many of these approaches remain theoretical, lacking practical integration with machine learning techniques, highlighting the need for more adaptive and implementable ethical frameworks in AI [32]. Our work represents a foundational effort to develop a unified framework that addresses ethical principles in AI systems, with a specific focus on granular levels of explainability and fairness. It builds upon existing approaches [33, 26] in specialized domains, extending them to tackle the unique challenges posed by modern AI ethics. In doing so, this paper introduces a novel direction for formalizing and verifying the ethical principles of AI systems.
@@
While the proposed framework is broadly applicable to the ethics verification of autonomous systems, this work specifically focuses on its application to AI. The scope of this work is to provide a conceptual foundation and framework for formalizing AI ethics using deontic and temporal logic. Rather than focusing on individual actions or decisions, our approach emphasizes system-level specifications. It involves defining ethical properties that AI systems should ideally meet. For example, an AI system that uses gender as a feature should avoid making decisions explicitly based on it. These properties ensure that systems are designed to identify and mitigate biases and ethical violations effectively. Defining properties like "forbidden to consider sensitive features in predictions" provides a way to analyze commonly discussed properties of AI in a unified manner. This abstraction also helps to reduce the difficulty in formalizing each action or each type of user [31]. Additional details on formalizing the properties of AI systems and their verification are provided in Section IV.
@@
The basic model for applying deontic logic to AI ethics uses first-order logic to define predicates (Table I) and axioms that capture ethical requirements. This model introduces variables such as $\mathit{x}$ to indicate an AI system, $\mathit{x}_{a}$ to indicate that system $\mathit{x}$ performs an action $\mathit{a}$, and predicates such as $\mathcal{E}(x)$ to indicate ethical behavior. Axioms 1.11.21.3, and 1.4 are then defined using these predicates to express ethical obligations, prohibitions, and permissions for AI systems. Building upon the basic model, an extended version incorporates temporal operators from temporal logic to reason about the ethical behavior of AI systems over time. This extension allows for the expression of more complex ethical requirements, such as the obligation for AI systems to maintain fairness over time or the prohibition of exhibiting bias. The temporal operators used in this model include “always” ($\Box$), “eventually” ($\diamond$), and “until” ($\mathcal{U}$) presented by Manna and Pnueli [34]. In this framework, $\mathit{x}$ persists across states, while actions cause state transitions. If no action influences the transition, the passage of time follows a default evolution independent of AI choices.
@@
Theorems III.3 and III.4 in the basic model of deontic logic for AI ethics explore the relationships between ethical obligations, prohibitions, and permissions for AI systems. These theorems employ first-order logic and the defined predicates to derive conclusions about the ethical behavior of AI systems. The proofs of these theorems rely on techniques such as modus ponens, contraposition, and proof by contradiction to establish the logical connections between the axioms and the derived statements. The general flavor of the theorems is to provide a rigorous foundation for reasoning about the ethical requirements of AI systems, demonstrating how the Axioms 1.11.21.3, and 1.4 can be used to infer specific obligations, prohibitions, and permissions in various contexts. By establishing these logical relationships, the theorems contribute to a comprehensive framework for analyzing and ensuring the ethical behavior of AI systems. Similar to this, the Theorems III.6 to III.15 in the extended model, which incorporates temporal logic operators, explore the ethical behavior of AI systems over time. These theorems focus on capturing the temporal aspects of ethical requirements, such as the obligation to maintain fairness or the prohibition of exhibiting bias. The proofs of these theorems utilize the semantics of the temporal operators, such as $\Box$(always), $\diamond$(eventually), and $\mathcal{U}$(until), in conjunction with the Axiom lists 2 to 3 and predicates defined (Table I) in the basic model. The general flavor of the theorems in the extended model is to provide a more expressive and nuanced framework for reasoning about the ethical behavior of AI systems, considering the dynamic and evolving nature of these systems. The theorems establish logical connections between the temporal properties of AI systems and their ethical obligations, allowing for the analysis of more complex and realistic scenarios. By incorporating temporal aspects, the extended model enables a deeper understanding of the long-term ethical implications of AI systems and provides a foundation for designing and verifying AI systems that behave ethically over time.
@@
The importance of this work lies in formalizing the axioms to define the ethical requirements of an AI system and its potential to provide a formal and verifiable framework for ensuring the ethical behavior of AI systems. Our experimental findings show the effectiveness of this formalization in assessing the ethics of real-world AI systems—Loan prediction and COMPAS. We evaluated the ethical aspects of the systems, wherein we defined specific properties that these systems must adhere to to be deemed ethical. Our results revealed that certain properties were indeed satisfied by the system, while others were not (Table III). The results demonstrated that applying deontic logic and temporal operators to AI ethics represents a significant step forward in formally specifying and verifying the ethical behavior of AI systems.
@@ II Related Works
The earlier work introduces the GenEth ethical dilemma analyzer [37], which utilizes inductive logic programming to infer principles for ethical actions. Dominance Act Utilitarianism (DAU), a deontic logic of agency, is another framework for encoding and analyzing obligations in autonomous systems. DAU frameworks are efficient in addressing safety-critical behaviors, such as adherence to traffic laws and avoidance of reckless actions [38]. Such frameworks can formalize ethical obligations in systems like self-driving cars, enabling systematic reasoning about social and moral responsibilities [39]. Additionally, several works employ the Belief-Desire-Intention (BDI) framework to formalize reasoning about moral agents [40, 41, 30]. This structure supports transparency and formal verification in ethical decision-making processes for robots.
Further, the literature explores the use of high-level action languages and Answer Set Programming to design ethical autonomous agents [42, 43, 44]. There are several works that propose using deontic logic to constrain robot behavior in ethically sensitive environments, as this type of logic helps interpret natural language directly [45, 26, 27]. These frameworks are also used for ethical reasoning in the healthcare domain, emphasizing accountability and transparency [46]. Additionally, various works focus on using deontic logic-based frameworks for formalizing ethical reasoning in AI systems [47, 48, 49].
@@
Thus, the literature suggests that rule-based ethical theories, particularly deontology, are essential for developing trustworthy AI systems [53]. However, considering the dynamic nature of AI, especially regarding fairness and explainability at a granular level, significant gaps remain. Our work addresses these gaps by introducing fairness and explainability at multiple granularities, including stable, transient, inherent, and retrofitted/artificial dimensions. These distinctions capture the evolving nature of AI and its complex decision-making processes, providing a more comprehensive approach to ethical verification. Furthermore, existing frameworks often overlook the impact of personal biases introduced during training and lack mechanisms to mitigate them effectively. To address this, we propose an iterative learning approach designed to identify and reduce the influence of personal biases in the system. While prior research highlights the gap between theoretical ethical reasoning and its practical application in autonomous agents [32], our framework bridges this divide. By implementing and testing the framework in real-world AI systems such as COMPAS and Loan prediction systems, we validate its effectiveness and ensure its applicability.
@@ III-A Preliminaries
Standard Deontic Logic (SDL) and Temporal Deontic Logic (TDL) represent two distinct variations within deontic logic. We incorporate both SDL and its extension, TDL, into our work. SDL formulas include classical propositional logic and it operates as a monadic deontic logic, meaning its operators (obligation, permission, forbidden) apply to individual formulas ($\varphi$); they are read as "it is obligatory that $\varphi$", "it is permissible that $\varphi$", and "it is forbidden that $\varphi$" respectively. Furthermore, they are cross-definable. For instance, $P\varphi:=\neg O(\neg\varphi)$, and $F\varphi:=O\neg\varphi$ This logical statement explains that permission ($P$) or forbidden ($F$) can be represented in terms of obligation ($O$). Temporal Deontic Logic expands SDL by integrating temporal aspects into norms and obligations, introducing operators such as $always(\Box),eventaully(\diamond),next$, and $until(\mathcal{U})$. For instance, $next\ \varphi$ means that the proposition $\varphi$ holds in the next time step. Similarly, $\varphi\mathcal{U}\psi$ means $\varphi$ is true until $\psi$ becomes true. We use the semantics of the combined logic as an extension of the Kripke-style possible world semantics of deontic logic with temporal operators, as described in reference [54]. While branching-time logic is often used to reflect future uncertainty, we adopt Linear Temporal Logic (LTL) due to its simplicity and relevance to AI verification, where obligations typically unfold along a single execution path. Moreover, LTL allows us to capture obligations or constraints that must persist, eventually hold, or be updated dynamically as the system learns or acts in a sequence of decisions. The alternativeness relation in our model preserves sequential consistency rather than enabling arbitrary time jumps. We recommend interested readers to refer [55, 54] for further details on the foundational principles.
In this section, we focus on formalizing the overall ethical behavior of an AI system. For the formalization, we use the predicates as shown in Table I. The predicate is a function that takes an input and returns a truth value. The following set of axioms has been formulated to articulate the necessary and sufficient characteristics for an AI system to be considered ethical. When modeling AI systems that require quantification over agents, actions, or time points, it is necessary to extend deontic and temporal logic with first-order logic formulas and quantifiers [56]. This framework serves as a foundational starting point for developing such an extended logical approach in this direction.
@@
Given a set $p$ of atomic propositions, the temporal deontic logic, TDL is defined as,
@@
We define the semantics of Temporal Deontic Logic (TDL) based on the foundational concepts presented in the work by [57, 54].
Definition III.2 (TDL Semantics).
Let $\mathcal{P}$ be a set of atomic propositions. A model for TDL is a tuple:
$\mathcal{M}=(S,R_{T},R_{O},D,I)$
where:
  • $S$ is a non-empty set of states,
  • $R_{T}\subseteq S\times S$ is the temporal accessibility relation,
  • $R_{O}\subseteq S\times S$ is the deontic accessibility relation,
  • $D$ is a constant domain of individuals,
  • $I$ is an interpretation function such that:
    • $I(p)\subseteq S$ for each $p\in\mathcal{P}$,
    • $P^{I}_{s}\subseteq D^{n}$ for each $n$-ary predicate $P$ at state $s$.
Let $\sigma$ be a variable assignment $\sigma:\text{Var}\to D$. The satisfaction relation $\mathcal{M},s,\sigma\models\phi$ is defined inductively as:
$\mathcal{M},s,\sigma\models p$ $\iff p\in I(s)$
$\mathcal{M},s,\sigma\models\neg\phi$ $\iff\mathcal{M},s,\sigma\not\models\phi$
$\mathcal{M},s,\sigma\models\phi\lor\psi$ $\iff\mathcal{M},s,\sigma\models\phi$ or $\mathcal{M},s,\sigma\models\psi$
$\mathcal{M},s,\sigma\models\phi\land\psi$ $\iff\mathcal{M},s,\sigma\models\phi$ and $\mathcal{M},s,\sigma\models\psi$
$\mathcal{M},s,\sigma\models\phi\to\psi$ $\iff\mathcal{M},s,\sigma\not\models\phi$ or $\mathcal{M},s,\sigma\models\psi$
$\mathcal{M},s,\sigma\models\Box\phi$ $\iff\forall s^{\prime}(sR_{T}s^{\prime}\Rightarrow\mathcal{M},s^{\prime},\sigma\models\phi)$
$\mathcal{M},s,\sigma\models\Diamond\phi$ $\iff\exists s^{\prime}(sR_{T}s^{\prime}\land\mathcal{M},s^{\prime},\sigma\models\phi)$
$\mathcal{M},s,\sigma\models\phi\mathcal{U}\psi$ $\iff\exists s^{\prime}(sR_{T}^{*}s^{\prime}\land\mathcal{M},s^{\prime},\sigma\models\psi$
and $\forall s^{\prime\prime}$ between $s$ and $s^{\prime}$:
$\mathcal{M},s^{\prime\prime},\sigma\models\phi)$
$\mathcal{M},s,\sigma\models O\phi$ $\iff\forall s^{\prime}(sR_{O}s^{\prime}\Rightarrow\mathcal{M},s^{\prime},\sigma\models\phi)$
$\mathcal{M},s,\sigma\models P\phi$ $\iff\exists s^{\prime}(sR_{O}s^{\prime}\land\mathcal{M},s^{\prime},\sigma\models\phi)$
$\mathcal{M},s,\sigma\models\neg P\phi$ $\iff\forall s^{\prime}(sR_{O}s^{\prime}\Rightarrow\mathcal{M},s^{\prime},\sigma\not\models\phi)$
$\mathcal{M},s,\sigma\models\forall v.\phi$ $\iff\forall d\in D,\ \mathcal{M},s,\sigma[v\mapsto d]\models\phi$
$\mathcal{M},s,\sigma\models\exists v.\phi$ $\iff\exists d\in D,\ \mathcal{M},s,\sigma[v\mapsto d]\models\phi$
@@
    If an AI system $\mathit{x}$ is ethical, then it is obligatory that $\mathit{a}$ is an ethical action and that $\mathit{x}$ performs $\mathit{a}$: $\forall x\forall a\ (\mathcal{E}(x)\rightarrow\mathcal{O}(\mathit{x_{a}}\land\mathcal{E}(\mathit{a})))$.
@@
    An ethical AI system $x$ is forbidden to perform an unethical action $\mathit{a}$: $\forall x\forall a\ (\mathcal{E}(x)\rightarrow\neg P(\mathit{x_{a}}\land\neg\mathcal{E}(a)))$.
@@
    For an ethical AI system $x$, performing an ethical action $\mathit{a}$ is permitted: $\forall x\forall a\ (\mathcal{E}(x)\rightarrow P(\mathit{x_{a}}\land\mathcal{E}(\mathit{a})))$.
@@
    An AI system following ethical guidelines performs ethical actions: $\mathcal{G}(x)\rightarrow\forall a\ \mathcal{E}(a)$
@@
$\mathcal{F}(\mathit{x})_{train}$ $x$ exhibits fairness in the training sample
$\mathcal{F}(\mathit{x})_{deploy}$ $x$ exhibits fairness during deployment
Based on the foundational principles outlined above and the domain knowledge, we have formulated Theorems  III.3 and  III.4 to ensure that the AI system adheres to ethical standards in all relevant dimensions. Theorem III.3 asserts that for an AI system to be ethical, it must follow all the ethical guidelines. It specifies that the system should not only be accurate in predicting outcomes but also do so in a manner that upholds principles such as fairness, transparency, and non-discrimination.
Theorem III.3.
@@
Assume $O(\mathit{x_{a}}\land\mathcal{E}(\mathit{a}))$. We want to prove $O(\mathit{x_{a}}\land\mathcal{G}(\mathit{x}))$.
  1. 1.
    From Axiom 1.4, $\mathcal{G}(\mathit{x})\rightarrow\mathcal{E}(\mathit{a})$
  2. 2.
    If the AI system performs action $\mathit{a}$ while following ethical guidelines, then action $\mathit{a}$ is ethical:
    $\mathit{x_{a}}\land\mathcal{G}(\mathit{x})\rightarrow\mathit{x_{a}}\land\mathcal{E}(\mathit{a})$
  3. 3.
    By Obligation on implication, since $\mathit{x_{a}}\land\mathcal{G}(\mathit{x})\rightarrow\mathit{x_{a}}\land\mathcal{E}(\mathit{a})$ is logically valid,
    $O(\mathit{x_{a}}\land\mathcal{G}(\mathit{x})\rightarrow\mathit{x_{a}}\land\mathcal{E}(\mathit{a}))$
  4. 4.
    Using the principle of obligation strengthening, and given $O(\mathit{x_{a}}\land\mathcal{E}(\mathit{a}))$ from our assumption (step 1), it follows that:
    $O(\mathit{x_{a}}\land\mathcal{G}(\mathit{x}))$
  5. 5.
    From steps 1 and 5, we conclude:
    $O(\mathit{x_{a}}\land\mathcal{E}(\mathit{a}))\rightarrow O(\mathit{x_{a}}\land\mathcal{G}(\mathit{x}))$
Theorem III.4 states that an AI system cannot be in a state where it is required to do something but not allowed to do it. This maintains logical and ethical consistency. Such a condition is fundamental in designing AI systems that reason ethically, as it ensures they are never blocked from doing what is morally required.
Theorem III.4.
@@
While Theorems  III.3 and  III.4 contribute to formalizing the general ethics of an AI system, it remains essential to develop rigorous formalizations for each ethical principle.